# HENet: Hybrid Encoding for End-to-end Multi-task 3D Perception from Multi-view Cameras

### v1

@@ 3.3 Independent BEV Feature Encoding
where $\mathbf{F}\in\mathbb{R}^{X\times Y\times C}$ is the BEV features, $\mathbf{W}$ denotes linear transform matrix, $f_{\text{avg}}$ indicates the global average pooling, and $\sigma$ represents the Sigmoid function. For the BEV encoder, we adopt three residual blocks and a simple FPN [22] following existing BEV-based methods [18, 11] to perform local feature integration on the BEV feature map. Notably, the adaptive feature selection and BEV encoders for different tasks share the same architecture but have different weights.
@@ 3.4 Decoders and Losses
where $\mathcal{L}_{depth}$ is the depth estimation loss and $\alpha$ is the loss weight.
@@ 4.1 Implementation Details
We train HENet in end-to-end manner for multi-tasks, including 3D object detection and BEV semantic segmentation in the same way as LSS [33]. We choose VovNet-99 [15, 31] with $896\times 1600$ image resolution for the large image encoder and select ResNet-50 [6] with $256\times 704$ image resolution for the small image encoder, respectively. For the input temporal sequence, we set short-term frame number $k=2$ and long-term frame number $n=9$. The weights of the hybrid image encoding network are initialized from pre-trained 3D detectors. As analyzed in Sec.4.4, we choose BEV grid sizes of 0.4m (256$\times$256 BEV size) and 0.8m (128$\times$128 BEV size) for 3D object detection and BEV semantic segmentation, respectively. The end-to-end multi-task models are trained for 60 epochs without CBGS. Besides, to further compare HENet with some single-task methods, we train the single 3D object detection models of HENet for 12 epochs with CBGS.
@@ 4.5 Single Task Results
For online evaluation, Table 3 shows HENet achieves state-of-the-art 3D object detection results. It is worth noting that HENet is built upon BEVDepth [18] and BEVStereo [17]. We can further improve the performance of HENet by incorporating stronger detectors, such as SparseBEV [26] and StreamPETR [37].
Table 3: Comparison of 3D object detection results on nuScences test set. The best and second best results are marked in red and blue.
@@ 4.6 Ablation
Table 5: Ablation of Hybrid Image Encoding Network. ‘Time’ denotes the overall training time. ‘BS’ indicates the batch size. FPS is measured on a single RTX3090 with FP32. Training cost (GPU memory and time) is estimated on 8$\times$×Tesla A800 GPUs with FP32.
@@
Hybrid Image Encoding Network. To demonstrate the effectiveness of the proposed hybrid image encoding network, we compare HENet with the two baseline methods and their ensemble model. As shown in Table 5, by combining with BEVDepth4D [18] and BEVStereo [17] through hybrid image encoding, HENet can significantly improve the 3D object detection performance. Compared to increasing resolution (model C), Hybrid Image Encoding Network can achieve higher accuracy with faster inference speed and lower training costs. Compared to increasing the frame number (model F), Hybrid Image Encoding Network can achieve higher accuracy with lower training costs. Notably, ensembling the results of the two baselines decreases the overall performance since the weaker BEVDepth4D [18] introduces many false positive detection results.

### v2

@@ 3.3 Independent BEV Feature Encoding
where $\mathbf{F}\in\mathbb{R}^{X\times Y\times C}$ is the BEV features, $\mathbf{W}$ denotes linear transform matrix, $f_{\text{avg}}$ indicates the global average pooling, and $\sigma$ represents the Sigmoid function. For the BEV encoder, we adopt three ResNet [6] residual blocks and a simple FPN [22] following existing BEV-based methods [18, 11] to perform local feature integration on the BEV feature map. Notably, the adaptive feature selection and BEV encoders for different tasks share the same architecture but have different weights.
@@ 3.4 Decoders and Losses
where $\mathcal{L}_{depth}$ is a binary cross entropy loss for depth estimation and $\alpha$ is the loss weight.
@@ 4.1 Implementation Details
We train HENet in end-to-end manner for multi-tasks, including 3D object detection and BEV semantic segmentation in the same way as LSS [33]. We choose VovNet-99 [15, 31] with $896\times 1600$ image resolution for the large image encoder and select ResNet-50 [6] with $256\times 704$ image resolution for the small image encoder, respectively. For the input temporal sequence, we set short-term frame number $k=2$ and long-term frame number $n=9$. The weights of the hybrid image encoding network are initialized from pre-trained 3D detectors. As analyzed in Sec.4.4, we choose BEV grid sizes of 0.4m (256$\times$256 BEV size) and 0.8m (128$\times$128 BEV size) for 3D object detection and BEV semantic segmentation, respectively. The end-to-end multi-task models are trained for 60 epochs without CBGS. Besides, to further compare HENet with some single-task methods, we train the single 3D object detection models of HENet for 12 epochs with CBGS [51].
@@ 4.5 Single Task Results
For online evaluation, Table 3 shows HENet achieves state-of-the-art 3D object detection results. To further demonstrate the effectiveness of HENet, we choose SparseBEV [26] as the baseline and incorporate it with HENet. We successfully implemented Hybrid Image Encoding of short-term ViT-L and long-term V2-99, proving the effectiveness of the Hybrid Image Encoding Network in reducing training costs. We achieve new state-of-the-art 3D detection results with 70.7 NDS and 64.5 mAP on nuScenes leaderboard (detection task/vision track).
Table 3: Comparison of 3D object detection results on nuScences test set. The best and second best results are marked in red and blue. †uses future frames. ‡uses test-time augmentation.
@@
BEVFormerV2 [44] InternImage-XL 64.8 58.0 0.448 0.262 0.342 0.238 0.128
BEVDet-Gamma [12] †‡ Swin-B 66.4 58.6 0.375 0.243 0.377 0.174 0.123
SparseBEV [26] V2-99 67.5 60.3 0.425 0.239 0.311 0.172 0.116
StreamPETR-Large [37] ViT-L 67.6 62.0 0.470 0.241 0.258 0.236 0.134
Hop [53] ViT-L 68.5 62.4 0.367 0.249 0.353 0.171 0.131
Far3D [13] ViT-L 68.7 63.5 0.432 0.237 0.278 0.227 0.130
HENet_Sp (Ours) † ViT-L & V2-99 70.7 64.5 0.402 0.235 0.237 0.155 0.129
@@ 4.6 Ablation
Table 5: Ablation of Hybrid Image Encoding Network on Detection Task. ‘Time’ denotes the overall training time. ‘BS’ indicates the batch size. FPS is measured on a single RTX3090 with FP32. Training cost (GPU memory and time) is estimated on 8$\times$×Tesla A800 GPUs with FP32.
@@
Table 6: Ablation of Hybrid Image Encoding Network on Multi-task.
Model Backbone Input Frames NDS mIoU
A BEVDepth4D R50 256$\times$704 9 frames 53.0 51.5
B BEVStereo V2-99 640$\times$1152 2 frames 58.0 55.8
C BEVStereo V2-99 896$\times$1600 2 frames 58.9 56.7
A+B HENet w/o AFFM V2-99 & R50 640$\times$1152 & 256$\times$704 2 & 7 frames 59.0 56.9
A+B HENet w/ AFFM V2-99 & R50 640$\times$1152 & 256$\times$704 2 & 7 frames 59.9 58.0
@@
Hybrid Image Encoding Network. To demonstrate the effectiveness of the proposed hybrid image encoding network, we compare HENet with the two baseline methods and their ensemble model. As shown in Table 5, by combining with BEVDepth4D [18] and BEVStereo [17] through hybrid image encoding, HENet can significantly improve the 3D object detection performance. Compared to increasing resolution (model C), Hybrid Image Encoding Network can achieve higher accuracy with faster inference speed and lower training costs. Compared to increasing the frame number (model F), Hybrid Image Encoding Network can achieve higher accuracy with lower training costs. Notably, ensembling the results of the two baselines by NMS decreases the overall performance since the weaker BEVDepth4D [18] introduces many false positive detection results.
For better comparison, we provide a multi-task ablation study, as shown in Table 6.
@@
Table 9: Multi-task results under different loss weights.
$\alpha_{depth}$ $\alpha_{cls}$ $\alpha_{bbox}$ $\alpha_{segveh}$ $\alpha_{segarea}$ $\alpha_{segdiv}$ NDS mIoU mIoUv mIoUa mIoUd
1 2.5 6 100 100 100 60.0 57.3 47.5 81.4 43.1
1 2.5 6 180 45 90 59.7 57.6 49.1 81.1 42.6
1 2.5 6 300 70 200 59.9 58.0 49.5 81.3 43.4
Loss Weights. Table 9 shows how we adjust the loss weight to obtain the results in Table 1. Specifically, we follow BEVDepth [18] to set $\alpha_{depth}=1$, $\alpha_{cls}=1.5$, and $\alpha_{box}=6$. We find that a larger segmentation loss weight achieves higher segmentation performance. Thus, we adjust the loss balance of each segmentation sub-task to obtain better overall segmentation results. The results show that carefully tuning loss weights can provide some improvement.